# Deflanderization for Game Dialogue: Balancing Character Authenticity with Task Execution in LLM-based NPCs

### v2

@@ 1 Introduction
First, user persona-LLMs are designed as purpose-built assistants that adapt to individual users’ preferences, backgrounds, and behavioral patterns  (Salemi et al., 2023; li2402personalized). These personalization systems leverage user-specific information to provide tailored responses, recommendations, and interactions. For example, LaMP (Large Language Models Meet Personalization) introduces comprehensive benchmarks for evaluating personalized text generation  (Salemi et al., 2023), while another work explores personalized dialogue agents that maintain consistent user preferences across conversations  (Zhang et al., 2018).
@@
These developments showcase the remarkable ability of modern LLMs to facilitate and embody given personas, with applications spanning from personalized user assistance to sophisticated multi-agent collaborations in virtual environments  (wu2023large; Jiang et al., 2022).
@@
Despite the rapid growth of LLM research, the entertainment field has remained relatively underexplored, particularly in traditional entertainment media creation such as video games. Conventional game development relies heavily on programmed logic, where in-game events and character interactions follow predetermined scripts and dialogue trees. To enhance player immersion and narrative depth, developers have begun incorporating LLMs as integral components of NPCs. This integration enables them to exhibit human-like behaviors and engage in dynamic, contextually-aware conversations with players  (Song et al., 2024).
@@
Recent advances in LLM-driven NPCs demonstrate significant potential for transforming player experiences. Cross-platform dialogue systems allow NPCs to maintain consistent interactions across both game environments and social platforms like Discord  (Song et al., 2024), creating unprecedented continuity in player-character relationships. Collaborative quest completion systems in Minecraft showcase how LLM-driven NPCs can work alongside human players to accomplish shared objectives  (Rao et al., 2024), while function-calling capabilities enable AI Game Masters or NPCs to manage complex game mechanics and narrative progression  (song2024you). Furthermore, specialized datasets like MCPDial  (alavi2024mcpdial) and PeaCoK  (Gao et al., 2023) provide rich persona-driven dialogue collections that enhance the authenticity and depth of NPC interactions, supporting the development of more sophisticated conversational agents in gaming environments.

### v3

@@ 1 Introduction
First, user persona-LLMs are designed as purpose-built assistants that adapt to individual users’ preferences, backgrounds, and behavioral patterns  (Salemi et al., 2023). These personalization systems leverage user-specific information to provide tailored responses, recommendations, and interactions. For example, LaMP (Large Language Models Meet Personalization) introduces comprehensive benchmarks for evaluating personalized text generation  (Salemi et al., 2023), while another work explores personalized dialogue agents that maintain consistent user preferences across conversations  (Zhang et al., 2018).
@@
These developments showcase the remarkable ability of modern LLMs to facilitate and embody given personas, with applications spanning from personalized user assistance to sophisticated multi-agent collaborations in virtual environments  (Jiang et al., 2022).
@@
Despite the rapid growth of LLM research, the entertainment field has remained relatively underexplored, particularly in traditional entertainment media creation such as video games. Conventional game development relies heavily on programmed logic, where in-game events and character interactions follow predetermined scripts and dialogue trees. To enhance player immersion and narrative depth, developers have begun incorporating LLMs as integral components of NPCs. This integration enables them to exhibit human-like behaviors and engage in dynamic, contextually-aware conversations with players  (Song et al., 2024b).
@@
Recent advances in LLM-driven NPCs demonstrate significant potential for transforming player experiences. Cross-platform dialogue systems allow NPCs to maintain consistent interactions across both game environments and social platforms like Discord  (Song et al., 2024b), creating unprecedented continuity in player-character relationships. Collaborative quest completion systems in Minecraft showcase how LLM-driven NPCs can work alongside human players to accomplish shared objectives  (Rao et al., 2024), while function-calling capabilities enable AI Game Masters or NPCs to manage complex game mechanics and narrative progression Song et al. (2024a). Furthermore, specialized datasets like MCPDial  Alavi et al. (2024) and PeaCoK  (Gao et al., 2023) provide rich persona-driven dialogue collections that enhance the authenticity and depth of NPC interactions, supporting the development of more sophisticated conversational agents in gaming environments.